How do I solve z^2(1-z^2)=16?
Answer by stanbon(75887)

solve z^2(1-z^2)=16
---
Let z^2 = x
----
x(1-x) = 16
x-x^2 = 16
x^2-x+16 = 0
x = [1 +- sqrt(1 -4*16)]/2
---
x = [1 +- sqrt(-63)]/2
---
x = [1 +- isqrt(63)]/2
----
z^2 = [1 +- isqrt(63)]/2
-----------------------------
etc.
Cheers,
Stan H.